OpenStudy (anonymous):

If ABCD congruent to QRST, mA = x – 10, and mQ = 2x – 30, what is the mA?

OpenStudy (anonymous):

x-10 = 2x-30 -10 = x - 30 20 = x

OpenStudy (anonymous):

mA = 20

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

If ABC congruent DEF, mA = 50, and mE = 30, what is mC? can you help with this

OpenStudy (anonymous):

actually the answer is 10, i just took the quick check.